OpenStudy (anonymous):

hi im lazy but please Find the slope of the line whose equation is x - 6y + 12 = 0. a. 1/6 b. -6 c. -1/6

OpenStudy (anonymous):

Slope is always in front of X y so simplify the equation then you will have your answer

OpenStudy (anonymous):

a

OpenStudy (anonymous):

1. Change x-6y+12=0 to -6y=-x-12 2. Multiply by -1 to get 6y=x+12 3. Divide by 6 to get y=(1/6)x+12 4. The coefficient of x is 1/6, and that's your slope

OpenStudy (anonymous):

thank you :)))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))

OpenStudy (anonymous):

Anytime